OpenStudy (anonymous):

f(x)= 2/sqrt of 3-x find the domain

OpenStudy (anonymous):

so fx=2/3-x

OpenStudy (anonymous):

\[\frac{ 2 }{ \sqrt{3-x} }\]

OpenStudy (anonymous):

opposite of division is multiply 2*3=6

OpenStudy (anonymous):

right

OpenStudy (anonymous):

3-6=-3 so x=-3

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

if so fan me and medal pls thanks

OpenStudy (anonymous):

you are supposed to find the domain

OpenStudy (anonymous):

domain of what? fx

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i figured it out though

OpenStudy (anonymous):

see this right here is how my teacher showed me and i'm showing u she said to me that's how u find the domain I don't know if u were showed another way

OpenStudy (anonymous):

the correct answer is {x|x<3} or (-infinite,3)

OpenStudy (anonymous):

i'm sorry I just screen shot this and showing my teacher

OpenStudy (anonymous):

You cannot square root a negative number and you cannot divide by zero, therefore x cannot equal 3 and cannot be greater than 3. So, x has to be less than 3. The (-infinity, 3) indicates the domain on a graph if you were to plot it. Notice that it is in round brackets instead of square brackets meaning that those values are not included.